OpenStudy (anonymous):

what is 120ft/day= in/min

OpenStudy (anonymous):

1 inch/minute

OpenStudy (a_clan):

120/(24*60) ft/min

OpenStudy (anonymous):

120 feet = 120*12 inches = 1440 inches 1 day = 24*60 minutes = 1440 minutes 1440 inches/ 1440 minutes = 1 inch/minute